OpenStudy (anonymous):

use logarithmic differentiation to find dy/dx for y=13(x^ln(x))

OpenStudy (anonymous):

take the ln of both sides, ln y = ln ( 13 x^ln x ) , and then take derivative of both sides so left side is y'/y , and right side, first add logs ln 13 + ln (x ^ln x) = ln 13 + ln x * ln x

OpenStudy (anonymous):

y'/y = 0 + 2ln x *1/x

OpenStudy (anonymous):

y' = 2ln x *1/x * y = 2 ln x * 1/x * 13 ( x ^ln x )

OpenStudy (anonymous):

thanks